OpenStudy (anonymous):

integral of:

OpenStudy (anonymous):

\[\int\limits_{?}^{?}(x+2)^2/(x^2+1)\]

OpenStudy (anonymous):

i think you have to multply out

OpenStudy (anonymous):

So integral of (x^2+4x+4)/(x^2+1) ? I tried that but idk what to do from there

OpenStudy (anonymous):

i would try \[\frac{x^2+1+4x+3}{x^2+1}=1+\frac{4x+3}{x^2+1}\] maybe that will help

OpenStudy (anonymous):

i suck at these so maybe there is a snappier way you are going to have to divide for sure, because as you see you get a 1 there the next part i am less sure about

OpenStudy (anonymous):

Can't I solve the integral of (4x+3)/(x^2+1) using that rule I forgot... Ax+B = ...

OpenStudy (anonymous):

partial fractions*

OpenStudy (anonymous):

no \(x^2+1\) does not factor, no partial fractions for this one

OpenStudy (anonymous):

break apart again to get a log and a arctangent i think

OpenStudy (anonymous):

What do you mean by break apart?

OpenStudy (anonymous):

ok you good to here \[\frac{x^2+1+4x+3}{x^2+1}=1+\frac{4x+3}{x^2+1}\]

OpenStudy (anonymous):

Yup I am

OpenStudy (anonymous):

then \[1+\frac{4x}{x^2+1}+\frac{3}{x^2+1}\]

OpenStudy (anonymous):

ohhh that's really smart. I see how to solve it now

OpenStudy (anonymous):

first term gives \(x\) last term is arctangent middle one is a simple u - sub

OpenStudy (anonymous):

oh ok said too much good luck

OpenStudy (anonymous):

Thanks so much for your help!

OpenStudy (anonymous):

yw